OpenStudy (anonymous):

solve 8sin^2(80)-2sqrt(3)*sin(40)-2cos(40)

OpenStudy (whpalmer4):

It's not an equation, so there's nothing to solve. You could evaluate or simplify it. \[8\sin^2(80) - 2\sqrt{3}*\sin(40) - 2\cos(40)\] You might try using the identity \[\sin(2x) = 2\sin(x)\cos(x)\] to convert that \(\sin^2(80)\) term into \(\sin(40)\) and \(\cos(40)\) terms.